# TableEval: A Real-World Benchmark for Complex, Multilingual, and Multi-Structured Table Question Answering

### v2

@@ Abstract
LLMs have shown impressive progress in natural language processing. However, they still face significant challenges in TableQA, where real-world complexities such as diverse table structures, multilingual data, and domain-specific reasoning are crucial. Existing TableQA benchmarks are often limited by their focus on simple flat tables and suffer from data leakage. Furthermore, most benchmarks are monolingual and fail to capture the cross-lingual and cross-domain variability in practical applications. To address these limitations, we introduce TableEval, a new benchmark designed to evaluate LLMs on realistic TableQA tasks. Specifically, TableEval includes tables with various structures (such as concise, hierarchical, and nested tables) collected from four domains (including government, finance, academia, and industry reports). Besides, TableEval features cross-lingual scenarios with tables in Simplified Chinese, Traditional Chinese, and English. To minimize the risk of data leakage, we collect all data from recent real-world documents. Considering that existing TableQA metrics fail to capture semantic accuracy, we further propose SEAT, a new evaluation framework that assesses the alignment between model responses and reference answers at the sub-question level. Experimental results have shown that SEAT achieves high agreement with human judgment. Extensive experiments on TableEval reveal critical gaps in the ability of state-of-the-art LLMs to handle these complex, real-world TableQA tasks, offering insights for future improvements. We make our dataset available here: https://github.com/wenge-research/TableEval.
@@
Junnan Zhu2thanks: Equal Contribution., Jingyi Wang1,211footnotemark: 1, Bohan Yu1,3, Xiaoyu Wu1, Junbo Li1, Lei Wang1,2, Nan Xu1,2thanks: Corresponding Author. 1 Beijing Wenge Technology Co., Ltd. 2 State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, CAS, Beijing, China 3 School of Advanced Interdisciplinary Sciences, University of Chinese Academy of Sciences, China junnan.zhu@nlpr.ia.ac.cn, nan.xu@wenge.comn
@@ 1 Introduction
We systematically assess LLMs on TableEval. Our extensive experiments reveal several notable observations: (1) Closed-source models (e.g., o1-preview, Claude-3.5-Sonnet) consistently lead in performance, but large open-source models (e.g., DeepSeek-R1, QwQ-32B-Preview) show significant promise with proper scaling and enhanced reasoning capabilities. (2) Understanding complex table structures remains a fundamental challenge, with performance drops of 10–15% on nested/hierarchical tables compared to flat tables. (3) Domain-specific and language-specific gaps persist, as models often struggle more in Chinese or specialized domains than in English or general settings.
@@ 2 Related Work
TableQA Methods. Question answering over tabular data has attracted increasing interest due to its potential to make complex data more accessible to non-experts. Early studies on tableQA primarily explored semantic parsing methods, treating the task as Text-to-SQL, where natural language queries and table schemas are directly converted into executable SQL Pasupat and Liang (2015); Wang et al. (2020); Shi et al. (2020); Guo et al. (2021). While these methods show great potential, they require a large amount of SQL-labeled training data and often assume that column headers are enough for reasoning, ignoring the valuable information in table cells. To overcome these problems, recent studies have shifted towards the retrieve-and-reasoning paradigm Eisenschlos et al. (2021); Zhong et al. (2022); Kumar et al. (2023); Lei et al. (2023) instead of directly generating SQL. Large-scale table pre-training Yin et al. (2020); Herzig et al. (2020); Liu et al. (2022); Jiang et al. (2022); Zhao et al. (2022) has further enhanced this shift, which has demonstrated strong performance on standard benchmarks. Furthermore, the advent of large language models enhanced with few-shot prompting Zhang et al. (2024a, b); Wang et al. (2024); Sui et al. (2024); Li et al. (2024) has significantly extended these capabilities, enabling more flexible, scalable, and generalizable reasoning over tabular data.
@@ 3.1 Dataset Construction
After collecting PDFs (and a limited number of HTML pages), we use parsing tools designed explicitly for PDF and HTML formats to extract tables into Excel worksheets, yielding 29,241 tables. To keep critical contextual information, we apply heuristic rules (see Figure 14 and Algorithm 1 in Appendix for details) to capture any adjacent text snippets near the tables, such as table captions or explanatory notes, so that the semantic context essential for question answering remains. Subsequently, three graduate students with finance and statistics backgrounds review the Excel files to correct parsing errors and ensure the table structures align accurately with the original PDFs. Each table is converted to Markdown format after reviewing. We provide an experiment on the impact of table format in Appendix C.3.
@@
We use K-means clustering for sampling and deduplication to enhance question diversity and ensure a balanced task distribution. Specifically, we use BGE-M3 Chen et al. (2024) to vectorize the questions for each task and cluster them by similarity, retaining only one representative query per cluster. This process results in 34,161 distinct questions. Of these, 2,325 (spanning 16 tasks) are candidates for the final dataset.
@@
3.2 Dataset Statistics
To address the need for more fine-grained evaluation of cross-domain capabilities, we have annotated each sample with detailed domain-specific subtask labels. For financial reports, we include subtasks such as financial metric retrieval, ratio and indicator trend analysis, and cross-period comparisons. Industry research reports encompass subtasks like industry performance comparison and financial indicator ranking. Academic papers contain experimental result queries and definition lookups. Government data includes price change calculations and industry growth rate queries.
Figure 3: Overview of Our SEAT evaluation method.
4 Evaluation Method for TableQA
@@ V H M
Table 2: Table Structure-wise Statistics.
@@
As shown in Figure 3, SEAT follows a structured process to evaluate the accuracy and relevance of model-generated responses, guided by the original question and a corresponding JSON-formatted reference answer. Each question is evaluated individually to provide a detailed assessment of the LLM’s performance. SEAT consists of two steps: (1) Extract key answers to each sub-question from the model’s final response, compare each with the reference answers to identify matching or divergent elements, and label each extracted answer as“true” or “false” based on its alignment with the reference. (2) Present the evaluation results in a JSON format, including the original question, the model’s response, and the evaluation record for each answer to facilitate tracking and verification. Compute the overall F1 for each QA pair by averaging the F1 scores across all sub-questions. For multi-turn conversational QA, first compute the average F1 for each turn, then average across all turns to obtain the final F1 score for the QA session.
Specifically, given a model-generated answer $A$ and a reference answer $R$, SEAT first converts them into structured forms using a Text2Json operation with a specific prompt (Figure 12) based on an LLM, which can be formalized as:
$J_{A}=\textrm{Text2Json}(A),J_{R}=\textrm{Text2Json}(R)$ (1)
where $J_{A}$ and $J_{R}$ are JSON formats that capture relevant answer components. If the reference answer is not already structured, SEAT applies Text2Json to extract a structured form.
Given a question $Q$ and the set of its sub-questions $S=\{S_{1},S_{2},...,S_{n}\}$. For each sub-question $S_{i}$, we verify whether $J_{A}$ correctly reflects the corresponding reference answer $r_{i}$ using LLM-as-a-judge Zheng et al. (2023a):
$\text{Match}(s_{i})=\begin{cases}1,&\text{if }\exists a_{j}\in J_{A}\text{, }\text{LLM-as-a-judge}(a_{j})=r_{i}\\ 0,&\text{otherwise}\end{cases}$ (2)
The results are recorded as a list, from which precision, recall, and F1 score will be calculated.
@@
5 Experimental Results
Effectiveness of SEAT. To assess the effectiveness of SEAT, we analyze the correlation between SEAT and human scores. Annotators rate model responses from 1 to 10 based on reference quality. We collect ratings for 5% randomly selected samples and evaluate the correlation using three metrics: 1) Pearson correlation coefficient ($r$), 2) Spearman rank coefficient ($\rho$), and 3) Kendall rank coefficient ($\tau$). As shown in Table 3, SEAT consistently achieves higher correlations with human judgments than existing metrics (F1 and ROUGE-L) and the LLM-as-a-judge. When using GPT-4o as the backbone, SEAT attains the highest agreement with human judgments. This improvement over the direct rating approach (LLM-as-a-judge) remains significant across all tested backbones, including Qwen-2.5-72B and Llama-3.3-70B, demonstrating SEAT’s robustness.
Main Results. Table 4 reports how well each LLM handles six main tasks in TableEval: Information Retrieval, Numerical Analysis, Reasoning, Data Analysis, Multi-turn Conversation, and Table Structure Understanding, along with the average score. Larger models generally perform better, with closed-source models like Claude-3.5-Sonnet, GPT-4o, and o1-Preview often leading the rankings. o1-Preview achieves the highest average score (83.43%), while Claude-3.5-Sonnet performs comparably (83.32%). Both perform well in complex tasks such as reasoning and data analysis, which require deep reasoning and domain knowledge. Among open-source models, DeepSeek-R1 shows strong, complex reasoning abilities and performs best in reasoning, numerical analysis, and table structure understanding. Qwen2.5-14B-Instruct significantly improves over its 7B version, demonstrating the benefits of increased parameters. Scaling up to 32B further enhances overall performance, with models like QwQ-32B-Preview and Qwen2.5-32B-Instruct improving across most metrics. However, even these large models still struggle with table structure understanding, suggesting that specialized training is necessary for tasks requiring layout comprehension. The significant gap between top closed-source LLMs and mid-sized open-source ones underscores the advantages that closed-source models may have, potentially through specialized training strategies, proprietary datasets, or more computational resources.
@@ 5 Experimental Results
Language-aspect Performance. Table 5 depicts the performance of different LLMs across tables in Simplified Chinese, Traditional Chinese, and English. Closed-source models demonstrate strong multilingual capabilities, maintaining high and consistent accuracy across these languages. Claude-3.5-Sonnet and o1-preview consistently achieve the top performance. A notable trend is a model’s development background’s influence on language strengths. Models trained in large Chinese corpora, such as the Qwen series and DeepSeek-R1, tend to perform best in English, followed by Simplified Chinese, with Traditional Chinese ranking lowest. In contrast, models primarily trained in English often achieve their highest accuracy in English or Traditional Chinese, while Simplified Chinese tends to perform worse.
@@
Key Insights and Observations. (1) Closed-Source Models Lead, but Open-Source Models Show Promise. Closed-source models (e.g., Claude-3.5-Sonnet, GPT-4o, o1-preview) consistently achieve top-tier performance across tasks, table structures, languages, and domains, benefiting from specialized training resources. However, large-scale open-source models (e.g., DeepSeek-R1, QwQ-32B-Preview) can be highly competitive when properly scaled and fine-tuned, indicating that open models have the potential to reduce the gap on complex tableQA with continued improvements. (2) Table Structure Understanding Remains a Challenge. Both closed- and open-source models struggle with complex table structures, particularly nested and hierarchical structures. Performance drops of 10–15% are observed compared to simple structures (flat tables), highlighting the importance of future research on explicit table structures or structure-aware encoding methods. (3) Domain and Language-Specific Performance Gaps. Open-source models like DeepSeek-R1 perform strongly in specific domains, such as government data, while closed-source models demonstrate strong cross-domain adaptability. Similarly, language-aspect evaluation shows that while models often perform well in English, their performance declines in Simplified or Traditional Chinese. These findings suggest that domain-specific of language-specific training strategies should be explored.
@@ Appendix A Dataset Construction Details
A.1 Task Formulation
A.2 Multi-answer Consistency Checking
@@
Information Retrieval Simple Lookup 369
@@
Numerical Analysis Sorting 223
@@
Data Analysis Causal Analysis 73
@@
Table Structure Understanding Table Size Detection 153

### v3

@@ Abstract
LLMs have shown impressive progress in natural language processing. However, they still face significant challenges in TableQA, where real-world complexities such as diverse table structures, multilingual data, and domain-specific reasoning are crucial. Existing TableQA benchmarks are often limited by their focus on simple flat tables and suffer from data leakage. Furthermore, most benchmarks are monolingual and fail to capture the cross-lingual and cross-domain variability in practical applications. To address these limitations, we introduce TableEval, a new benchmark designed to evaluate LLMs on realistic TableQA tasks. Specifically, TableEval includes tables with various structures (such as concise, hierarchical, and nested tables) collected from four domains (including government, finance, academia, and industry reports). Additionally, TableEval features cross-lingual scenarios with tables in Simplified Chinese, Traditional Chinese, and English. To reduce potential data leakage, we curate data from recent real-world documents. Considering that existing TableQA metrics fail to capture semantic accuracy, we further propose SEAT, a new evaluation framework that assesses the alignment between model responses and reference answers at the sub-question level. Experimental results have shown that SEAT achieves high agreement with human judgment. Extensive experiments on TableEval reveal critical gaps in the ability of state-of-the-art LLMs to handle these complex, real-world TableQA tasks, offering insights for future improvements. We make our dataset available here: https://github.com/wenge-research/TableEval.
@@
Junnan Zhu2thanks: Equal Contribution., Jingyi Wang1,211footnotemark: 1, Bohan Yu1,3, Xiaoyu Wu1, Junbo Li1, Lei Wang1,2, Nan Xu1,2thanks: Corresponding Author. 1 Beijing Wenge Technology Co., Ltd. 2 State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, CAS, Beijing, China 3 School of Advanced Interdisciplinary Sciences, University of Chinese Academy of Sciences, China junnan.zhu@nlpr.ia.ac.cn, nan.xu@wenge.com
@@ 1 Introduction
We systematically assess LLMs on TableEval. Our extensive experiments reveal several notable observations: (1) Closed-source models (e.g., o1-preview, Claude 3.5 Sonnet) consistently lead in performance, but large open-source models (e.g., DeepSeek-R1, QwQ-32B-Preview) show significant promise with proper scaling and enhanced reasoning capabilities. (2) Understanding complex table structures remains a fundamental challenge, with performance drops of 10–15% on nested/hierarchical tables compared to flat tables. (3) Domain-specific and language-specific gaps persist, as models often struggle more in Chinese or specialized domains than in English or general settings.
@@ 2 Related Work
TableQA Methods. Question answering over tabular data has attracted increasing interest due to its potential to make complex data more accessible to non-experts. Early studies on TableQA primarily explored semantic parsing methods, treating the task as Text-to-SQL, where natural language queries and table schemas are directly converted into executable SQL Pasupat and Liang (2015); Wang et al. (2020); Shi et al. (2020); Guo et al. (2021). While these methods show great potential, they require a large amount of SQL-labeled training data and often assume that column headers are enough for reasoning, ignoring the valuable information in table cells. To overcome these problems, recent studies have shifted towards the retrieve-and-reasoning paradigm Eisenschlos et al. (2021); Zhong et al. (2022); Kumar et al. (2023); Lei et al. (2023) instead of directly generating SQL. Large-scale table pre-training Yin et al. (2020); Herzig et al. (2020); Liu et al. (2022); Jiang et al. (2022); Zhao et al. (2022) has further enhanced this shift, which has demonstrated strong performance on standard benchmarks. Furthermore, the advent of LLMs enhanced with few-shot prompting Zhang et al. (2024a, b); Wang et al. (2024); Sui et al. (2024); Li et al. (2024) has significantly extended these capabilities, enabling more flexible, scalable, and generalizable reasoning over tabular data.
@@ 3.2 Dataset Construction
After collecting PDFs (and a limited number of HTML pages), we use parsing tools designed explicitly for PDF and HTML formats to extract tables into Excel worksheets, yielding 29,241 tables. To keep critical contextual information, we apply heuristic rules (see Figure 14 and Algorithm 1 in Appendix for details) to capture any adjacent text snippets near the tables, such as table captions or explanatory notes, so that the semantic context essential for question answering remains. Subsequently, three graduate students with finance and statistics backgrounds review the Excel files to correct parsing errors and ensure the table structures align accurately with the original PDFs. Each table is converted to Markdown format after reviewing. We provide an experiment on the impact of table format in Appendix B.3.
@@
We use K-means clustering within each task for sampling and deduplication to enhance question diversity and ensure a balanced task distribution. For each task, we set the expected number of QA pairs as the number of clusters. Specifically, we use BGE-M3 Chen et al. (2024) to obtain question embeddings and cluster them by similarity, retaining only one representative query per cluster. This process yields 34,161 distinct QA pairs. Among these, 5,422 pairs (spanning 16 tasks) are selected as candidates for the final test set, with the remainder serving as the candidate training set.
@@ V H M
Table 2: Table Structure-wise Statistics. V, H, M, Hier, Conc, and Nested represent vertical, horizontal, matrix, hierarchical, concise, and nested table types.
Figure 3: Overview of Our SEAT evaluation method.
Second-stage annotation employs a structured answer extraction process. We use LLMs to determine if the original question needs to be decomposed into sub-questions and to extract JSON-formatted answers from the annotated responses. This process filters redundant and low-relevance information from natural language responses, refining the answer content to improve comparison accuracy. The resulting JSON-formatted answers then undergo two rounds of human verification and one round of expert review to ensure accuracy and consistency with the original table data and question context. The comprehensive annotation process averaged approximately \$2.5 per sample.
3.3 Dataset Statistics
Our dataset provides comprehensive coverage of various domain-specific subtasks, supporting a more fine-grained evaluation of cross-domain capabilities. For financial reports, we include subtasks such as financial metric retrieval, ratio and indicator trend analysis, and cross-period comparisons. Industry research reports encompass subtasks like industry performance comparison and financial indicator ranking. Academic papers contain experimental result queries and definition lookups. Government data includes price change calculations and industry growth rate queries.
4 Evaluation Method for TableQA
@@
As shown in Figure 3, SEAT follows a structured process to evaluate the accuracy and relevance of model-generated responses, guided by the original question and a corresponding JSON-formatted reference answer. SEAT consists of two steps: (1) Extract key answers to each sub-question from the model’s final response, compare each with the reference answers to identify matching or divergent elements, and label each extracted answer as“true” or “false” based on its alignment with the reference. (2) Present the evaluation results in a JSON format, including the original question, the model’s response, and the evaluation record for each answer to facilitate tracking and verification. Compute the overall F1 for each QA pair by averaging the F1 scores across all sub-questions. For multi-turn conversational QA, first compute the average F1 for each turn, then average across all turns to obtain the final F1 score for the QA session.
Specifically, given a model-generated answer $A$ and a reference answer $R$, both in natural language, SEAT converts them into structured forms and evaluates their alignment in a single step. It uses an LLM guided by a specific prompt (see Figure 12), which can be formalized as:
$\text{J}_{\text{eval}}=\text{Text2Json}(A,R)$ (1)
where Text2Json instructs the LLM to parse both $A$ and $R$. The LLM is prompted to first extract key answer components from $A$ corresponding to the question, then compare these components against the information in $R$ to determine their correctness (labeling each as "true" or "false"). The final output, $\text{J}_{\text{eval}}$, is a single JSON object that contains the structured representations of both answers alongside the correctness judgments for each sub-question. Precision, recall, and the F1 score are then calculated from this structured output. This integrated approach ensures that the semantic comparison is directly tied to the structuring process, providing a robust and consistent evaluation.
5 Experiment
5.1 Experimental Setup
We evaluate 19 models, ranging from 7B to 671B parameters, to assess their performance on TableEval. The evaluated models include open-source and closed-source models to comprehensively compare different model architectures and capabilities. For open-source models, we evaluate Qwen2.5-Instruct series Qwen et al. (2025): Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct; Qwen2.5-Coder-Instruct-32B Hui et al. (2024); QwQ-32B-Preview; DeepSeek series DeepSeek-AI et al. (2024): DeepSeek-V2-Lite-Chat (16B), DeepSeek-Coder-V2-Lite-Instruct (16B), DeepSeek-V2.5-1210 (236B), DeepSeek-V3 (671B), DeepSeek-R1 (671B) DeepSeek-AI et al. (2025); glm-4-9b-Chat (9B) GLM et al. (2024); Llama-3 series Grattafiori et al. (2024): Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct. For closed-source models, we evaluate qwen-max-2024-09-19, Claude 3.5 sonnet-20241022, GPT-4o-mini-2024-07-18, GPT-4o-2024-11-20 OpenAI et al. (2024), and o1-preview. We set the temperature to 0 and use a fixed seed (42) to make the model’s output more deterministic. For models that do not allow the configuration of parameters (e.g., temperature or seed), we use their default settings.
@@ 5.1 Experimental Setup
5.2 Effectiveness of SEAT
To assess the effectiveness of SEAT, we analyze the correlation between SEAT and human scores. Annotators rate model responses from 1 to 10 based on reference quality. We collect ratings for 5% randomly selected samples and evaluate the correlation using three metrics: 1) Pearson correlation coefficient ($r$), 2) Spearman rank coefficient ($\rho$), and 3) Kendall rank coefficient ($\tau$). As shown in Table 3, SEAT consistently achieves higher correlations with human judgments than existing metrics (F1 and ROUGE-L) and the LLM-as-a-judge. When using GPT-4o as the backbone, SEAT attains the highest agreement with human judgments. This improvement over the direct rating approach (LLM-as-a-judge) remains significant across all tested backbones, including Qwen-2.5-72B and Llama-3.3-70B, demonstrating SEAT’s robustness.
5.3 Main Results
Table 4 reports how well each LLM handles six main tasks in TableEval: Information Retrieval, Numerical Analysis, Reasoning, Data Analysis, Multi-turn Conversation, and Table Structure Understanding, along with the average score. Larger models generally perform better, with closed-source models like Claude 3.5 Sonnet, GPT-4o, and o1-preview often leading the rankings. o1-preview achieves the highest average score (83.43%), while Claude 3.5 Sonnet performs comparably (83.32%). Both perform well in complex tasks such as reasoning and data analysis, which require deep reasoning and domain knowledge. Among open-source models, DeepSeek-R1 shows strong, complex reasoning abilities and performs best in reasoning, numerical analysis, and table structure understanding. Qwen2.5-14B-Instruct significantly improves over its 7B version, demonstrating the benefits of increased parameters. Scaling up to 32B further enhances overall performance, with models like QwQ-32B-Preview and Qwen2.5-32B-Instruct improving across most metrics. However, even these large models still struggle with table structure understanding, suggesting that specialized training is necessary for tasks requiring layout comprehension. The significant gap between top closed-source LLMs and mid-sized open-source ones underscores the advantages that closed-source models may have, potentially through specialized training strategies, proprietary datasets, or more computational resources.
@@ 5.3 Main Results
Language-aspect Performance. Table 5 depicts the performance of different LLMs across tables in Simplified Chinese, Traditional Chinese, and English. Closed-source models demonstrate strong multilingual capabilities, maintaining high and consistent accuracy across these languages. Claude 3.5 Sonnet and o1-preview consistently achieve the top performance. A notable trend is a model’s development background’s influence on language strengths. Models trained in large Chinese corpora, such as the Qwen series and DeepSeek-R1, tend to perform best in English, followed by Simplified Chinese, with Traditional Chinese ranking lowest. In contrast, models primarily trained in English often achieve their highest accuracy in English or Traditional Chinese, while Simplified Chinese tends to perform worse.
@@
Key Insights and Observations. (1) Closed-Source Models Lead, but Open-Source Models Show Promise. Closed-source models (e.g., Claude 3.5 Sonnet, GPT-4o, o1-preview) consistently achieve top-tier performance across tasks, table structures, languages, and domains, benefiting from specialized training resources. However, large-scale open-source models (e.g., DeepSeek-R1, QwQ-32B-Preview) can be highly competitive with test-time scaling, indicating that open models have the potential to reduce the gap on complex TableQA with continued improvements. (2) Table Structure Understanding Remains a Challenge. Both closed- and open-source models struggle with complex table structures, particularly nested and hierarchical structures. Performance drops of 10–15% are observed compared to simple structures (flat tables), highlighting the importance of future research on explicit table structures or structure-aware encoding methods. (3) Domain and Language-Specific Performance Gaps. Open-source models like DeepSeek-R1 perform strongly in specific domains, such as government data, while closed-source models demonstrate strong cross-domain adaptability. Similarly, language-aspect evaluation shows that while models often perform well in English, their performance declines in Simplified or Traditional Chinese. These findings suggest that domain-specific of language-specific training strategies should be explored.
@@ 7 Limitations
Acknowledgements
This work is supported by Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park No.Z231100007423016, the National Natural Science Foundation of China under Grants No.62206287. We sincerely thank the anonymous reviewers for their insightful comments and constructive suggestions.
@@ Appendix A Dataset Construction Details
Information
Retrieval
Simple Lookup 369
@@
Numerical
Analysis
Sorting 223
@@
Data
Analysis
Causal Analysis 73
@@
Table Structure
Understanding
Table Size Detection 153
@@ A.1 Multi-answer Consistency Checking
A.2 Structured Answer Extraction
Appendix B Additional Experimental Analysis
B.1 Robustness of SEAT